OpenStudy (jadedry):

Given x = 2sin (nt + pi/3) and y= 4sin (nt + pi/6), Find the Cartesian equation of the locus of the point (x,y) as t varies.

OpenStudy (jadedry):

I've already expressed x and y in terms of sin nt and cos nt: \[X = \sin nt + \sqrt 3 \cos nt\] \[Y= 2\sqrt3 \sin nt + 2 \cos nt\] I'm not exactly sure as to how the question wants me to proceed. Thanks in advance!

OpenStudy (michele_laino):

hint: we can apply this identity: \[\begin{gathered} x = 2\sin \left( {nt + \frac{\pi }{3}} \right) = 2\sin \left( {nt + \frac{\pi }{6} + \frac{\pi }{6}} \right) = 2\sin \left( {\left( {nt + \frac{\pi }{6}} \right) + \frac{\pi }{6}} \right) = \hfill \\ \hfill \\ = 2\sin \left( {nt + \frac{\pi }{6}} \right)\cos \left( {\frac{\pi }{6}} \right) + 2\cos \left( {nt + \frac{\pi }{6}} \right)\sin \left( {\frac{\pi }{6}} \right) = \hfill \\ \hfill \\ = 2 \cdot \frac{y}{4} \cdot \cos \left( {\frac{\pi }{6}} \right) + 2 \cdot \sqrt {1 - {{\left( {\frac{y}{4}} \right)}^2}} \cdot \sin \left( {\frac{\pi }{6}} \right) \hfill \\ \end{gathered} \]

OpenStudy (jadedry):

Thanks for the hint! I'll look at some more videos and see if I can solve it!

OpenStudy (jadedry):

@Michele_Laino I'm sorry to bother you, but I'm not how to deal with this. My book says: \[4x^2 + y^2 - 2\sqrt3xy = 4\] I worked backwards from the answer, and it seems to simplify to: \[12\sin^2 + 12 \cos^2 = 4\] I don't know where the 4 comes from at all!

OpenStudy (michele_laino):

here is the next step: \[\begin{gathered} x = 2\sin \left( {nt + \frac{\pi }{3}} \right) = 2\sin \left( {nt + \frac{\pi }{6} + \frac{\pi }{6}} \right) = 2\sin \left( {\left( {nt + \frac{\pi }{6}} \right) + \frac{\pi }{6}} \right) = \hfill \\ \hfill \\ = 2\sin \left( {nt + \frac{\pi }{6}} \right)\cos \left( {\frac{\pi }{6}} \right) + 2\cos \left( {nt + \frac{\pi }{6}} \right)\sin \left( {\frac{\pi }{6}} \right) = \hfill \\ \hfill \\ = 2 \cdot \frac{y}{4} \cdot \cos \left( {\frac{\pi }{6}} \right) + 2 \cdot \sqrt {1 - {{\left( {\frac{y}{4}} \right)}^2}} \cdot \sin \left( {\frac{\pi }{6}} \right) = \hfill \\ \hfill \\ = \frac{y}{2} \cdot \frac{{\sqrt 3 }}{2} + \frac{{\sqrt {4 - {y^2}} }}{2} \hfill \\ \end{gathered} \] so we can write: \[x = \frac{y}{2} \cdot \frac{{\sqrt 3 }}{2} + \frac{{\sqrt {4 - {y^2}} }}{2}\]

OpenStudy (jadedry):

So: \[4x^2 = 2y^2 +2\sqrt (3y - y^3) +4 \]

OpenStudy (michele_laino):

sorry, I have made a typo, here is the right final step: \[x = \frac{y}{2} \cdot \frac{{\sqrt 3 }}{2} + \frac{{\sqrt {16 - {y^2}} }}{4}\]

OpenStudy (jadedry):

Ah! let me rework it then.

OpenStudy (michele_laino):

and we can rewrite such equation, like this: \[x - \frac{y}{2} \cdot \frac{{\sqrt 3 }}{2} = \frac{{\sqrt {16 - {y^2}} }}{4}\]

OpenStudy (michele_laino):

please square both sides

OpenStudy (jadedry):

I'm squaring it now, just bear with me please!

OpenStudy (michele_laino):

hint: \[{\left( {x - \frac{y}{2} \cdot \frac{{\sqrt 3 }}{2}} \right)^2} = {\left( {\frac{{\sqrt {16 - {y^2}} }}{4}} \right)^2}\]

OpenStudy (jadedry):

\[x^2 - \frac{ \sqrt3 xy }{ 2 } + \frac{ 3y^2 }{ 16 } = \frac{16 - y^2? }{ 16 } \]

OpenStudy (michele_laino):

correct!

OpenStudy (jadedry):

therefore: \[16x^2 - 8\sqrt3 xy +3y^2 = 16 - y^2 = 16x^2 - 8\sqrt3 xy + 4y^2 = 16\]

OpenStudy (jadedry):

\[4x^2 - 2\sqrt3 xy + y ^2 = 4\]

OpenStudy (jadedry):

Ah! thank you so much for your help. I see how my book came to the answer!

OpenStudy (michele_laino):

:)

OpenStudy (jadedry):

Alright, I'll be closing the question now. Thanks once again for your help and patience!

OpenStudy (michele_laino):

:)